# MultiCamCows2024 - A Multi-view Image Dataset for AI-driven Holstein-Friesian Cattle Re-Identification on a Working Farm

### v1

@@ Abstract
We present MultiCamCows2024, a farm-scale image dataset filmed across multiple cameras for the biometric identification of individual Holstein-Friesian cattle exploiting their unique black and white coat-patterns. Captured by three ceiling-mounted visual sensors covering adjacent barn areas over seven days on a working dairy farm, the dataset comprises $101,329$ images of $90$ cows, plus the underlying original CCTV footage. The dataset is provided alongside full computer vision recognition baselines, that is both a supervised and self-supervised learning framework for individual cow identification trained on cattle tracklets. We report a performance above $96\%$ single image identification accuracy from the dataset and demonstrate that combining data from multiple cameras during learning enhances self-supervised identification. We show that our framework enables fully automatic cattle identification, barring only the simple human verification of tracklet integrity during data collection. Crucially, our study highlights that multi-camera, supervised and self-supervised components in tandem not only deliver highly accurate individual cow identification but also achieve this efficiently with no labelling of cattle identities by humans at all. We argue that this improvement in efficacy has practical implications for livestock management, behaviour analysis, and agricultural monitoring. For full reproducibility and practical ease of use, we publish all key software and code including re-identification components and the species detector with this paper.
@@
Animal Biometrics , Smart Farming , Cattle , Holstein-Friesian , Computer Vision , Machine Learning , Contrastive Learning , Self-supervised Learning , Re-Identification , Imageomics , Artificial Intelligence
PACS:
$0$ , $1$
MSC:
$0$ , $1$
@@ 1 Introduction
Monitoring of Animals. Today, sensor-based approaches [1] for the study of animals play a crucial role in various aspects of biological and veterinary research and practice [2, 3, 4], and provide a critical tool for implementing animal conservation and welfare policies [5]. Vision-based Artificial Intelligence (AI) methodologies, in particular, have become the technology of choice for applications related to biodiversity assessment [6, 7, 8] and smart farming [9, 10]. In these settings modern machine-learning approaches often boost precision, efficiency, and flexibility of the monitoring tasks at hand. AI tasks range from species recognition, tracking and pose estimation, body condition scoring, behaviour detection and individual animal re-identification (Re-ID) [11].
Motivation for Dairy Cattle Identification. In agriculture specifically, keeping track of individuals across time and space by performing Re-ID tasks is one key requirement that underpins many applications ranging from welfare assessment to production management activities such as growth monitoring, estrus detection and precision insemination timing [12, 13, 14, 15]. We note that Holstein-Frisian cattle are a particularly important species to focus efforts on since they form a foundation for dairy farming [16] in many regions of the world [17]. Moreover, the utilisation of 24/7 monitoring methods for this species promises wide-reaching improvements in both efficiency and also welfare in the dairy sector [18]. Expanded roll-out of such domain-specific AI systems, however, first requires further improvements both in terms of a reduction in the amount of human input required for training, but also in the accuracy of system outputs. In any case, cattle monitoring via Computer Vision (CV) is a particularly promising avenue due to the remote operation of cameras without any need for physical sensor contact or reader stations positioned very close to the animals.
@@
Scenarios, Efficiency, and Transferability. Using both individual and species-encompassing knowledge, based on large datasets followed by limited fine-tuning has been acknowledged as one efficient approach for the training and running of farm surveillance applications. To this end, Gao et.al. [22], for instance, created a relatively large single-camera dataset to enable experimentation regarding the minimisation of human-labelling for the identification of individual Holstein-Friesian cows from top-down CCTV. However, using a single top-down camera limits identification to one view. It provides little knowledge regarding realistic monitoring performance for scenarios where multiple cameras are needed to cover a farm – possibly with overlapping fields of view.
@@ 2.1 Re-Identification and Farm Systems
Supervised Deep Learning for Static Domains. Across domains, fully-supervised deep learning has achieved remarkable results regarding target categorisation [25, 26] and localisation [27, 28, 29] and is the technology of choice for Re-Identification systems. Yet, its heavy reliance on fixed, annotated data at training time does not deal well with some common problems, such as the introduction of new animals with novel patterns. Training large classifiers from scratch under such settings is inefficient and often not feasible in practical, economically constrained settings.
Learning in Dynamic Data Domains. Adaptive transfer learning and self-supervised learning [30, 31, 32], for instance, have helped address such dataset-wide scenarios by transferring meaningful knowledge from an existing model onto unlabelled datasets that retain some structure (e.g. the same animal species – even if patterned completely differently). Reliant on a ResNet backbone and spectral embedding analysis, Luo et al. [11], showed that a straightforward framework for human Re-ID tasks can operate by utilising a single, carefully crafted embedding space. Able to propose accurate identification over smaller batches of data efficiently, Pathak et al. [33] refined this toolkit through inner fine-graining across spatial subregions of latent space. Therien et al. [34] extended this Re-ID framework for 3D depth sensor applications by utilising PointNet [35] and ViT  [36]. We note that these approaches are, in principle, fully applicable to the Holstein-Friesian cattle Re-ID task.
@@ 2.2 Species Recognition
CNN-based Object Recognition. Deep residual CNNs and their derivatives [40, 27, 41] are the most widespread tool-kits for semantic detection and segmentation. The introduction of region proposal networks (RPNs) [28] allowed for performing such tasks faster and perform real-time segmentation tasks. Later, pyramid-shaped networks utilised stacking region selection with inter-layer knowledge to acquire multi-level features [42]. Knowledge of the previously trained semantic segmentation model can also be further distilled for transfer learning over advanced and detailed features [29]. Based on these and related [43] approaches, cattle recognition and even behaviour analysis has been applied to Holstein-Friesians [44] including work by Andrew et.al. [24] proposing a drone-based, automatic CNN framework to monitor Holstein-Friesian cows from the air.
Transformer-based Object Recognition. Precise segmentation performance in higher-resolution images is often reduced when using a traditional CNN architecture due to the fixed-sized operations within blocks. An alternative, first used in language models [45], introduces various versions of Vision Transformers (ViT) to the task of object recognition [36]. However, ViTs and similar architectures face challenges regarding effective token sampling and correlation acquisition. In response, Fayyaz et.al. [46] boosted ViT performance with an adaptive token sampling module and Yu et.al. [47] replaced attention layers with a pooling module and built a lightweight model on top of a general MetaFormer architecture.
@@ 2.3 Individual Identification
Construction of Latent Spaces for Individuals. In order to classify a population, a network is usually tasked with mapping the visual appearance of individuals into a distinctly clustered latent space, where clusters relate to individuals. Contrastive learning can be used to shape such a space and metric learning [21], in particular, aims at building the space such that distances in latent space reflect the relationships between the instances of the population [49]. Chen et.al. [50] built SimCLR in which faster and more accurate results can be learned via a selection of prior data augmentation. To handle unsupervised learning tasks for open set problems, Oord et.al. [51] applied auto-regressive models to encode input datasets into higher dimensional spaces. We utilise metric learning for clustering the individually-indicative latent space building on work by Gao et al. [23]. Before discussing these experiments we will first outline how the core dataset was collected and prepared.
@@ 3.2 Tracklet Generation and Refinement
Species Detection and Tracklet Generation. We first applied species detection [52] to the farm footage in order to extract cattle locations for each frame in the CCTV video (see Fig. 3). We trained this species detector for cow torso recognition using pre-trained weights from ImageNet [53] for ResNet-152[41] as our backbone. We adapted the tools from NVIDIA Computer Vision Toolkit [^1] https://github.com/NVIDIA/retinanet-examples to perform head-oriented, rotated bounding box detections, which are later normalised so that all images to have the same orientation – i.e. the cow is always seen walking left-to-right in the dataset. Due to the wide visual range of views, we limited detection to a region of interest (RoI) in the cattle lane for each camera (see Fig. 1). Finally, we fine-tuned our species detector via Adam [54] using $5,606$ annotated frames across three selected views. Fig. 4 shows a training and performance overview for this species detector.
@@ 3.3 Cattle Re-Identification
Latent Space Construction. After generating uniformly-labelled tracklets, we proceeded to conduct supervised (Fig. 5(a)) and self-supervised (Fig. 5(b)) learning. Subsequently, we evaluate performance. Both identification approaches employed ImageNet pre-trained ResNet18 backbones responsible for embedding images into feature vectors of the latent space. Before inputting images into the models, we normalize them to have a uniform size and photometric variance. Our training dataset comprises tracklets from the initial five days of data. Day six was used for validation – and day seven as the test set. Training was conducted via the AdamW optimiser [55] at a learning rate of $0.001$ sampled at $201$ epochs (chosen due to minimal validation loss at this point) using an NVIDIA RTX2080Ti system with 48GB of GPU memory. Training ran for $2873$ minutes overall. The full training process is illustrated in Fig. 4 (left).
Loss Function, Evaluation, and Experimental Setup. Within the supervised framework, we employ a model incorporating cross-entropy loss and a one-hot classifier. To address variations in quantity amongst individuals (i.e. class imbalances), we sample images of individuals with equal probability for training. Within the self-supervised framework, a linear layer was utilised to generate multi-dimensional feature vectors for metric learning purposes. The NTXentLoss [50] was used as a contrastive loss function. To ensure that the ground-truth identities of the cows could not ‘leak’ into the self-supervision process, pseudo-labels were used (effectively randomizing the order in which cows appeared daily). For testing our unsupervised classifier, the ground-truth labels were used and KNN clustering [56] was applied to evaluate performance. As part of the comparative analysis, we also substituted the ImageNet weights from our frozen self-supervision counterpart into the supervised classifier to analyse the model’s ability to learn the intricate mapping between cow body feature maps and their corresponding labels.
@@ 4 Results
Re-Identification Performance. Here we present results on the task of individual identification based on training from tracklets as described above. Table 1 represents the quantitative performance of our supervised classifier and self-supervised identifier, respectively, where we compared the results of training for each independent camera and also with them all combined. For our classifier trained via supervised methods, the performance of merging the camera views reached the best accuracy result at $96.06\%$ with the cropped image size being $128\times 128$ using a $32$D-embedding. This is marginally inferior to camera $1$ performance ($96.78\%$), but is significantly improved compared to previously published single camera performance in slightly different settings benchmarked at $76.9\%$ accuracy as reported for Cows2021 [22]. Self-supervision demonstrated a small overall increase in performance when merging multiple views together as shown in Table 1. When training with our ResNet18 backbone with images resized to $128\times 128$, combining views leads to an accuracy of $96.36\%$, boosting the performance of camera $1$, $2$ and $3$ by $0.48\%$, $3.62\%$, and $39.67\%$ respectively. The relatively large gap between the performance of camera $3$ versus the combined views was due to the cows in camera $3$ being mostly stationary (they tend to stop walking after the milk-race, so group together making it difficult to disambiguate individuals accurately). Cattle detection in dense groups (due to a camouflaging effect of the coat pattern) is a truly difficult task not solved to the same quality levels as standard object detection. Fig 7 and A.2 provide visualisations of embeddings obtained during inference from self-supervision. Additionally, in Fig. 6 we show the confusion matrices of ground truth vs. prediction. Note that pattern similarity remains the major reason for misidentification. For instance, cow $035$ is often misclassified as cow $031$ due to the similarity of their body patterning from the rear and the side.
@@
$AllCameras$ 95.79 [95.70, 95.88] 93.92 [93.87, 93.96]
$Camera1@Hid32$ 96.06 [95.85, 96.28] 93.74 [93.65, 93.83]
$Camera2@Hid32$ 94.85 [94.61, 95.09] 91.02 [90.93, 91.11]
$Camera3@Hid32$ 91.92 [91.71, 92.12] 46.93 [46.71, 47.14]
$AllCameras$ 95.41 [95.31, 95.50] 93.61 [93.56, 93.66]
$Camera1@Hid64$ 96.30 [96.12, 96.50] 91.50 [91.42, 91.59]
$Camera2@Hid64$ 94.99 [94.77, 95.21] 90.68 [90.58, 90.78]
$Camera3@Hid64$ 93.37 [93.18, 93.57] 73.82 [73.66, 73.98]
Table 2: Confidence Interval Table. Shown are two hyper-parameter settings with Hid32 (upper rows) and Hid64 (lower rows). Subset Mean (Mean) and Confidence intervals (CI) of both supervised and self-supervised classifiers from single and multiple cameras. We sampled a randomised subset 25% the size of the original test dataset and ran inference 100 times.

### v2

# Holstein-Friesian Re-Identification using Multiple Cameras and Self-Supervision on a Working Farm

@@ Abstract
We present MultiCamCows2024, a farm-scale image dataset filmed across multiple cameras for the biometric identification of individual Holstein-Friesian cattle exploiting their unique black and white coat-patterns. Captured by three ceiling-mounted visual sensors covering adjacent barn areas over seven days on a working dairy farm, the dataset comprises $101,329$ images of $90$ cows, plus underlying original CCTV footage. The dataset is provided with full computer vision recognition baselines, that is both a supervised and self-supervised learning framework for individual cow identification trained on cattle tracklets. We report a performance above $96\%$ single image identification accuracy from the dataset and demonstrate that combining data from multiple cameras during learning enhances self-supervised identification. We show that our framework enables automatic cattle identification, barring only the simple human verification of tracklet integrity during data collection. Crucially, our study highlights that multi-camera, supervised and self-supervised components in tandem not only deliver highly accurate individual cow identification, but also achieve this efficiently with no labelling of cattle identities by humans. We argue that this improvement in efficacy has practical implications for livestock management, behaviour analysis, and agricultural monitoring. For reproducibility and practical ease of use, we publish all key software and code including re-identification components and the species detector with this paper, available at https://tinyurl.com/MultiCamCows2024.
@@
Animal Biometrics , Smart Farming , Holstein-Friesian , Self-supervised Learning , Re-Identification
@@ 1 Introduction
Monitoring of Animals. Today, sensor-based approaches for the study of animals including animal biometrics [1] play a crucial role in various aspects of biological and veterinary research and practice [2]. They provide a critical tool for the implementation of animal conservation and welfare policies [3, 4]. Vision-based Artificial Intelligence (AI) methodologies, in particular, have become the technology of choice [5] for applications related to species conservation and biodiversity assessment [6, 7] as well as smart farming [8, 9]. In these settings, modern machine learning approaches often boost precision, efficiency, and flexibility of the monitoring tasks at hand. AI tasks range from species recognition, tracking and pose estimation, body condition scoring, behaviour detection and individual animal re-identification (Re-ID).
Motivation for Dairy Cattle Identification. In agriculture specifically, keeping track of individuals in time and space by performing Re-ID tasks is a key requirement that underpins many applications ranging from welfare assessment to production management activities such as growth monitoring, estrus detection, and precision insemination timing. We note that Holstein-Frisian cattle are a particularly important species for which to focus efforts, as they form a foundation for dairy farming in many regions of the world. Moreover, the utilisation of 24/7 monitoring methods for this species promises wide-reaching improvements in both efficiency and also welfare in the dairy sector [10]. However, expanded roll-out of such domain-specific AI systems first requires further improvements both in terms of a reduction in the amount of human input required for training, but also in the accuracy of system outputs. In any case, self-supervised cattle monitoring via Computer Vision (CV) is a particularly promising avenue due to the remote operation of cameras without the need for physical sensor contact or reader stations positioned very close to the animals.
@@
Scenarios, Efficiency, and Transferability. Using both individual and species-encompassing knowledge, based on large datasets followed by limited fine-tuning has been acknowledged as one efficient approach for the training and running of farm surveillance applications. To this end, Gao et.al. [14], for instance, created a relatively large single-camera dataset to enable experimentation regarding the minimisation of human-labelling for the identification of individual Holstein-Friesian cows from top-down CCTV. Their work demonstrated the importance of self-supervision by reducing reliance on human labeling, establishing it as a strong backbone candidate for learning transferable representations to identify targets and guide future research. However, using a single top-down camera limits identification to one view. It provides little knowledge regarding realistic monitoring performance for scenarios where multiple cameras are needed to cover a farm – possibly with overlapping fields of view. Consequently, customised re-training becomes unavoidable when applying a single-camera model across a multi-camera setup with varying angles, heights, and hardware configurations, significantly increasing human labour as the number and diversity of cameras grow.
@@ 2.1 Re-Identification and Farm Systems
Supervised Deep Learning for Static Domains. Across domains, fully-supervised deep learning has achieved remarkable results regarding target categorisation and localisation, and is the technology of choice for Re-Identification systems. Yet, its heavy reliance on fixed, annotated data at training time does not deal well with some common problems, such as the introduction of new animals with novel patterns. Training large classifiers from scratch under such settings is inefficient and often not feasible in practical, economically constrained settings.
Learning in Dynamic Data Domains. Adaptive transfer learning and self-supervised learning have aided address such dataset-wide scenarios by transferring meaningful knowledge from an existing model onto unlabelled datasets that retain some structure (e.g. the same animal species – even if patterned completely differently). Reliant on a ResNet backbone and spectral embedding analysis, Able to propose accurate identification over smaller batches of data efficiently, Pathak et al. [16] refined earlier human Re-ID tools through inner fine-graining across spatial subregions of latent space. Therien et al. [17] then extended this Re-ID framework for 3D depth sensor applications by utilising PointNet and ViT. We note that these approaches are, in principle, fully applicable to the Holstein-Friesian cattle Re-ID task.
@@ 2.2 Species Recognition
CNN-based Object Recognition. Deep residual CNNs and their derivatives are the most widespread tool-kits for semantic detection and segmentation. The introduction of region proposal networks (RPNs) allowed for performing such tasks faster and perform real-time segmentation tasks. Later, pyramid-shaped networks utilised stacking region selection with inter-layer knowledge to acquire multi-level features. Knowledge of the previously trained semantic segmentation model can also be further distilled for transfer learning over advanced and detailed features. Based on these and related approaches, animal recognition and even behaviour analysis has not only been applied to wild species [21], but also to Holstein-Friesian cattle [22] including work by Andrew et.al. [12] proposing a drone-based, automatic CNN framework to monitor Holstein-Friesian cows from the air.
Transformer-based Object Recognition. Precise segmentation performance in higher-resolution images is often reduced when using a traditional CNN architecture due to the fixed-sized operations within blocks. An alternative, first used in language models, introduces various versions of Vision Transformers (ViT) to the task of object recognition. However, ViTs and similar architectures face challenges regarding effective token sampling and correlation acquisition. In response, Fayyaz et.al. [23] boosted ViT performance with an adaptive token sampling module and Yu et.al. [24] replaced attention layers with a pooling module and built a lightweight model on top of a general MetaFormer architecture.
@@ 2.3 Individual Identification
Construction of Latent Spaces for Individuals. In order to classify a population, a network is usually tasked with mapping the visual appearance of individuals into a distinctly clustered latent space, where clusters relate to individuals. Contrastive learning can be used to shape such a space and metric learning [13], in particular, aims at building the space such that distances in latent space reflect the relationships between the instances of the population. Chen et.al. [26] built SimCLR in which faster and more accurate results can be learned via a selection of prior data augmentation. They utilised Normalized Temperature-scaled Cross Entropy Loss (NTXentLoss) to maximise the distance between all unique labels. We utilise metric learning for clustering the individually-indicative latent space building on work by Gao et al. [15]. Before discussing these experiments we will first outline how the core dataset was collected and prepared.
@@ 3.2 Tracklet Generation and Refinement
Species Detection and Tracklet Generation. Adapted from previous works by Gao et.al.[14], we first applied species detection [13] to the farm footage in order to extract cattle locations for each frame in the CCTV video (see Fig. 3). We trained this species detector for cow torso recognition using pre-trained weights from ImageNet for ResNet-152 as our backbone. We adapted the tools from NVIDIA Computer Vision Toolkit [^1] https://github.com/NVIDIA/retinanet-examples to perform head-oriented, rotated bounding box detections, which are later normalised so that all images to have the same orientation – i.e. the cow is always seen walking left-to-right in the dataset. Due to the wide visual range of views, we limited detection to a region of interest (RoI) in the cattle lane for each camera (see Fig. 1). Finally, we fine-tuned our species detector via Adam optimiser using $5,606$ annotated frames across three selected views. Fig. 4 shows a training and performance overview for this species detector.
@@ 3.3 Cattle Re-Identification
Latent Space Construction. After generating uniformly-labelled tracklets which now reflect the 90 total cows we have collected features from, we proceeded to conduct supervised (Fig. 5-(a)) and self-supervised (Fig. 5-(b)) learning. Subsequently, we evaluate performance. Both identification approaches employed ImageNet pre-trained ResNet18 backbones responsible for embedding images into feature vectors of the latent space. Before inputting images into the models, we normalize them to have a uniform size and photometric variance. Our training dataset comprises tracklets from the initial five days of data. Day six was used for validation – and day seven as the test set. Training was conducted via the AdamW optimiser at a learning rate of $0.001$ sampled at $201$ epochs (chosen due to minimal validation loss at this point) using an NVIDIA RTX2080Ti system with 48GB of GPU memory. Training ran for $2873$ minutes overall. The full training process is illustrated in Fig. 4 (left).
Loss Function, Evaluation, and Experimental Setup. Within the supervised framework, we employ a model incorporating cross-entropy loss and a one-hot classifier on top of the reshaped images. To address variations in quantity amongst individuals (i.e. class imbalances), we sample images of individuals with equal probability for training. Within the self-supervised framework, a linear layer was utilised to generate multi-dimensional feature vectors for metric learning purposes. The NTXentLoss [26] was used as a contrastive loss function. It encourages distance reduction of items having the same label in latent space while moving these items away from those with contrasting labels. To ensure that the ground-truth identities of the cows could not ‘leak’ into the self-supervision process, pseudo-labels were used (effectively randomizing the order in which cows appeared daily). For testing our unsupervised classifier, the ground-truth labels were used and KNN clustering was applied to evaluate performance. As part of the comparative analysis, we also substituted the ImageNet weights from our frozen self-supervision counterpart into the supervised classifier to analyse the model’s ability to learn the intricate mapping between cow body feature maps and their corresponding labels. Both the supervised and self-supervised frameworks have been trained using NVIDIA GeForce RTX 2080 Ti GPUs.
@@ 4 Results
Re-Identification Performance. Here we present results on the task of individual identification based on training from tracklets as described above. Table 1 represents the quantitative performance of our supervised classifier and self-supervised identifier, respectively, where we compared the results of training for each independent camera and also with them all combined. For our classifier trained via supervised methods, the performance of merging the camera views reached the best accuracy result at $96.06\%$ with the cropped image size being $128\times 128$ using a $32$D-embedding. This is marginally inferior to camera $1$ performance ($96.78\%$), but is significantly improved compared to previously published single camera performance in slightly different settings benchmarked at $76.9\%$ accuracy as reported for Cows2021 [14]. Self-supervision demonstrated a small overall increase in performance when merging multiple views together as shown in Table 1. When training with our ResNet18 backbone with images resized to $128\times 128$, combining views leads to an accuracy of $96.36\%$, boosting the performance of camera $1$, $2$ and $3$ by $0.48\%$, $3.62\%$, and $39.67\%$ respectively. With more images causing dataset imbalance, the relatively large gap between the performance of camera $3$ versus the combined views was due to the cows in camera $3$ being mostly stationary (they tend to stop walking after the milk-race, so group together making it difficult to disambiguate individuals accurately). Cattle detection in dense groups (due to a camouflaging effect of the coat pattern) is a truly difficult task not solved to the same quality levels as standard object detection. Fig. 7 and A.2 provide visualisations of embeddings obtained during inference from self-supervision. Additionally, in Fig. 6 we show the confusion matrices of ground truth vs. prediction. Note that pattern similarity remains the major reason for misidentification. For instance, cow $035$ is often misclassified as cow $031$ due to the similarity of their body patterning from the rear and the side.
K-Fold Cross Validation. In addition, we applied K-Fold Cross Validation (KFCV) to fully evaluate model test performances in both fully- and self-supervised settings(Table 2). We chose the standard $k=10$ for cross validation after shuffling all the data for our fully-supervised pipeline. Since data formatting and training logic are significantly different for our self-supervised pipeline where cows were categorised in dates before in ID labels, we chose a customised $k=7$ based on 7 days from our data source for cross-validation. To avoid deviating from our training details, we sequentially cycle through each of 7 days as the test set and perform training and validation from the remaining 6 days, with 5 days as the training set and one as validation.
@@
$AllCameras$ $99.01\pm 1.09$ [98.18, 99.83] $91.70\pm 5.84$ [85.87, 97.53]
$Camera1@Hid32$ $99.76\pm 0.28$ [99.55, 99.97] $82.70\pm 8.60$ [74.11, 91.29]
$Camera2@Hid32$ $98.72\pm 0.95$ [98.00, 99.43] $84.36\pm 6.80$ [77.57, 91.15]
$Camera3@Hid32$ $98.45\pm 1.39$ [97.31, 99.58] $52.95\pm 13.37$ [39.60, 66.31]
$AllCameras$ $99.63\pm 0.22$ [99.47, 99.81] $87.87\pm 2.84$ [85.03, 90.71]
$Camera1@Hid64$ $99.91\pm 0.08$ [99.85, 99.98] $76.34\pm 9.60$ [66.75, 85.92]
$Camera2@Hid64$ $99.50\pm 0.32$ [99.26, 99.74] $83.37\pm 4.61$ [78.77, 87.98]
$Camera3@Hid64$ $99.32\pm 0.27$ [99.12, 99.53] $47.61\pm 10.66$ [36.96, 58.26]
Table 2: K-Fold Cross Validation Table. K-Fold Cross Validation of both fully- and self-supervised frameworks. Shown are two hyper-parameter settings with Hid32 (upper rows) and Hid64 (lower rows). Subset Mean (Mean) with Standard Deviation and Confidence Intervals (CI) from single and multiple cameras.